OpenStudy (anonymous):

How long it will take for $300 to double when invested at 6% annual interest compounded twice a year?

OpenStudy (campbell_st):

change the interest the half yearly r = 3% then 600 = 300(1 + 3/100)^n 2 = 1.03^n to find n you need logarithms ln(2) = nln(1.03) n = ln(2)/ln(1.03) n = 23.44 6 month periods. or its 11.72 years the alternative is to set up the equation in a calculator and keep change n until you get the desired value of 2

OpenStudy (anonymous):

@campbell_st Thank you very much